OpenStudy (anonymous):

write an equation for the quadratic function whose graph contains the vertex (3, 8) and point (5, -1)

OpenStudy (anonymous):

y = a(x - h)^2 + k -> -1 = a(5 - 3)^2 + 8 -> -1 = 4a + 8 -> a = -9/4 y = (-9/4)(x - 3)^2 + 8 is the vertex form

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

uw!